One most commonly used utility function is the Cobb-Douglas utility function which of the form 𝑼(𝑿, 𝒀) = 𝑿 πœΆπ’€ 𝜷 where Ξ± and Ξ² are positive constants. a. Show that this function exhibits diminishing marginal utility for both goods X and Y (4mks) b. Show that the indifference curves of this utility function are convex (i.e show that is there is diminishing marginal rate of substitution between X and Y)Β 


1
Expert's answer
2022-01-31T09:58:23-0500

If the slope of marginal utility ( derivative of marginal utility, is negative, consuming more units of a good increases utility by smaller and smaller increment shows the case of diminishing marginal utility.


"U=(X^{\alpha}Y^\beta)"

"Mu_x= \alpha X^{\alpha-1}Y^\beta"


"\alpha-1" Gives a negative figure since "\alpha" is less than 1


"\delta Mu_x=\frac{\delta Mu_x}{\delta X}= \alpha(\alpha-1)X^{\alpha-2}Y^\beta"


"=" "\frac{\alpha(\alpha-1)Y^\beta}{X^{\alpha-2}}"

Since "\alpha-1" is negative, then "\delta Mu_x" is negative indicating a diminishing marginal utility


"Mu_y= \beta X^\alpha Y^{\beta-1}"


"\beta-1" Gives a negative figure since "\beta" is less than 1


"\delta Mu_y=\frac{\delta Mu_y}{\delta y}"


"= \beta(\beta-1)X^\alpha Y^{\beta-2}"


"=" "\frac {\beta (\beta-1) X^\alpha} {Y^{\beta-2}}"

Since "\beta-1" is negative, then "\delta Mu_y" is negative indicating a diminishing marginal utility